In order to join a voice lesson class, you pay a P1,500.00 pesos fee, then P500.00 for

each class you go to. What is the average cost per class? Graph the function to show

whether it forms a straight line or a curve.​


1
Expert's answer
2021-10-07T15:11:06-0400

Solution: pesos fee = 1500

each class fee = 500

no of class = "x"

average fees(A) per class = "\frac{1500+500x}{x}"

if "x" = 1

average fees = "\frac{1500+500(1)}{1}" = 2000

if "x" = 2

average fees = "\frac{1500+500(2)}{2}" = 1250

if "x" = 3

average fees = "\frac{1500+500(3)}{3}" = 1000

if "x" = 4

average fees ="\frac{1500+500(4)}{4}" = 875

now we got number of classes and average fess , now we plot the graph between these points

"x" = number of classes and A = average fees

"x" = 1 , A = 2000

"x" = 2 , A = 1250

"x" = 3 , A = 1000

"x" = 4 , A = 875